You can move around with your keyboard
plus
minus
left
right
Drop pin on the map to start a new report
space
Press space again to adjust the location
space
Back to all reports

Light number 1 bossard court, lu7 1DF

Reported via mobile in the Street Lights category anonymously at 20:15, Sun 20 November 2022

Sent to Central Bedfordshire Council less than a minute later. Council ref: 378484.

I have reported this light not working for weeks now the actual light was replaced quite quickly but now nothing else seems to be being done, also the alleyway very close by also has street lights out and on dogget street maybe an electrical problem in the area.

RSS feed of updates to this problem Receive email when updates are left on this problem.

Updates

  • Thank you for contacting us with information relating to a Highways issue. Our engineers have inspected and found a fault with the power supply which is owned and maintained by the electricity provider. We have requested that they investigate and make a repair (their target timescale is 28 days). We will continue to monitor this issue until it is resolved. Please contact us again after that time if the light is still out.

    State changed to: In progress

    Posted by Central Bedfordshire Council at 09:03, Wed 23 November 2022

  • Numerous engineers have apparently tried to fix it, the light itself is new compared to the old orange light but I have never seen it working, it.s now not been working for about 2 months

    State changed to: Open

    Posted by Central Bedfordshire Council at 20:36, Sun 18 December 2022

  • Numerous engineers have apparently tried to fix it, the light itself is new compared to the old orange light but I have never seen it working, it’s now not been working for about 2 months

    Posted anonymously at 20:36, Sun 18 December 2022
    Still open, via questionnaire

  • Questionnaire filled in by problem reporter

    Posted by Central Bedfordshire Council at 22:14, Sunday 15 January 2023

  • Questionnaire filled in by problem reporter

    State changed to: Fixed

    Posted anonymously at 22:14, Sunday 15 January 2023

This report is now closed to updates. You can make a new report in the same location.